The balance sheet provides a snapshot of the financial condition of a company. Investors and analysts use the information given on the balance sheet and other financial statements to make several interpretations regarding the companys financial condition and performance Blue Hamster Manufacturing Inc. is a hypothetical company. Suppose it has the following balance sheet items reported at the end of its first year of operation. For the second year, some parts are still incomplete. Use the information given to complete the balance sheets for Blue Hamster Manufacturing Inc. for the years ending December 31, Year 2 and 1, respectively. Blue Hamster Manufacturing Inc. Balance Sheet For the Year ended December 31 Year 2 Year 1 Year 2 Year 1 Liabilities and equity Assets Current assets Cash and equivalents Accounts receivable Inventories Total current assets Net fixed assets: Net plant and equipment Current liabilities $0 $5,859 $33,203 $92,250 Accounts payable $33,750 Accruals $99,000 Notes payable $225,000 Total current liabilities $0 $0 $31,250 $31,250 $93,750 $125,000 $42,188 $123,750 $281,250 Long-term debt $117,188 $275,000 Total debt $156,250 Common equity: Common stock Retained earnings Total common equity Total liabilities and equity $304,688 $243,750 $%131,250 $375,000 $500,000 $468,750 Total assets $625,000 $500,000 $625,000Statement #1: Blue Hamsters accumulated owed financial obligations decreased from Year 1 to Year 2. This statement is , because : Long-term debt decreased from $33,203 at the end of Year 1 to $31,250 by the end of Year 2 Accruals actually increased from $0 in Year 1 to $5,859 at the end of Year 2 The accounts payable account remained $0 for both Years 1 and 2 Notes payable actually increased from $33,203 to $31,250 between Years 1 and 2. Statement #2: On December 31 of Year 2, Blue Hamster Manufacturing Inc. had $115,312 of actual money that it could have spent immediately This statement is because The funds recorded in Blue Hamsters accounts receivable account represents funds that are either cash or can be converted into cash almost immediately Blue Hamsters Year 2 cash and equivalents balance is $290,250. The funds recorded in Blue Hamsters cash and equivalents account represents f almost immediately. cash or can be converted into cash ScreenshotStatement #3: One way to interpret the change in Blue Hamsters accounts receivable balance from Year 1 to Year 2 is that more customers purchased new items on credit rather than paying off existing credit accounts. This statement is because: The decrease from $42,188 to $33,750 implies a net decrease in accounts receivable and that more customers are paying off their receivables balances than are buying on credit The change from $99,000 to $123,750 reflects a net accumulation of new credit sales. The $8,438 increase in accounts receivable means either that Year 1s existing credit customers are not paying off their owed balances and new or existing customers are making additional purchases on credit, or that Year 1s credit customers have repaid their owed balances and Year 2 credit sales have exceeded Year 1s credit sales. Based on your understanding of the different items reported in the balance sheet and the information they provide, which statement regarding Blue Hamster Manufacturing Inc.s balance sheet is consistent with U.S. Generally Accepted Accounting Principles (GAAP)? The companys assets should be listed from those carrying the largest balance to those with the smallest balance. The companys assets should be listed in alphabetical order The companys assets should be listed in the order in which they are to be converted into cash.
Answer #1

1) Statement 1 is False because

Accruals actually increased from $0 in year 1 to $5859 at the end of year 2

All other statements are false as can be verified from the given balance sheet

2) Statement 2 is True because

The funds recorded in Blue Hamster’s cash and equivalents account represents funds that are either cash or can be converted into cash almost immediately

Cash and cash equivalents in year 2 = total current assets in year 2 – accounts receivable in year 2 – inventory in year 2 = 281250 – 42188-123750 = $115,312

Thus Net Plant & equipment in year 2 = Total assets in year 2 – total current assets in year 2

= 625000-281250 = $343,750

3) Statement 3 is True because

Change in accounts receivable from year 1 to year 2 = accounts receivable in year 2 – accounts receivable in year 1 = 42,188-33,750 = $8438

The $8438 increase in accounts receivable means either that year 1’s existing credit customers are not paying off their owed balances and new or existing customers are making additional purchases on credit, or that year 1’s credit customers have repaid their owed balances and year 2 credit sales have exceeded Year 1’s credit sales.

4) As per U.S. GAAP and from the given balance sheet:

The company’s assets should be listed in the order in which they are to be converted into cash.

Total current liabilities in year 2 = accounts payable in year 2 + accruals in year 2 + Notes payable in year 2 = 0 + 5859 + 33203 = $39062

Retained earnings in year 2 = total common equity in year 2 – common stock in year 2 = 468750 – 304688 = $164,062